OpenStudy (anonymous):

If the measures of the angles of a quadrilateral ABCD are in the extended ratio m

OpenStudy (anonymous):

4x+5x+3x+6x=360 18x=360 x=20 each angle=4*20,5*20,6*20,3*20=80,100,120,60 

OpenStudy (anonymous):

awesome thank you.

OpenStudy (anonymous):

yw